# Spectral properties of symmetrized AMV operators

### v2

@@ 5. $L^{2}$ convergence
In this smooth context, the function $V(\cdot,r)$ is obviously continuous for any $r>0$. Since $M$ is compact, this implies that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$). Then Lemma 3.2 applies and ensures that $\tilde{\Delta}_{r}$ is a bounded self-adjoint operator acting on $L^{2}(M,\mathrm{vol}_{g})$. The compactness of $M$ also ensures that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is locally Ahlfors regular: there exists a constant $C>1$ such that for any $x\in M$ and $r\in(0,\mathrm{diam}(M)]$,
@@ 5.1. Convergence in the sense of distributions
Recall that $C_{m}$ is defined in (3). For any $x\in M\backslash\partial M$, we let $\exp_{x}$ be the exponential map centered at $x$. We identify $T_{x}M$ with $\mathbb{R}^{m}$ and write $\mathbb{B}^{m}_{r}(v)$ for the Euclidean ball in $\mathbb{R}^{m}$ centered at $v$ with radius $r>0$. Then there exists $\delta>0$ such that the restriction of $\exp_{x}$ to $\mathbb{B}^{m}_{\delta}(0)$ is a diffeomorphism onto its image; recall that the injectivity radius $\mathrm{i}_{M}(x)$ of $M$ at $x$ is the supremum of the set of such numbers $\delta$. We let $J_{x}$ be the Radon-Nikodym derivative of the measure $(\exp_{x}^{-1})_{\#}\mathrm{vol}_{g}$ with respect to the Lebesgue measure $\mathcal{L}^{m}$. It is well-known that for any $\xi\in\mathbb{B}_{\mathrm{i}_{M}}^{m}(0)$,
@@ 5.3. Uniform bound
We wish now to provide a uniform $L^{\infty}$ bound for the functions $-\tilde{\Delta}_{r}\psi$ where $r$ is in a neighborhood of zero.
@@
for some positive definite, symmetric $(m-1)$-square matrix $A_{(\xi,t)}$. From the non-degeneracy of the metric and a Lipschitz bound, we have that there exist $C,\tilde{c}>0$ such that for all $t,s\in[0,\epsilon],\xi\in[-3,3]^{m-1},v\in\mathbb{R}^{m}$ we have
@@
Consider $\tilde{c}>0$ given by (123). We know that there exists $M>0$ such that the map $\sqrt{\cdot}:[\tilde{c},+\infty)\rightarrow\mathbb{R}$ is $M$-Lipschitz. By homogeneity in $|v|$ of (125), we can assume that $|v|=1$. Then the Lipschitz condition and (124) yield
@@
Set $\mathbb{H}^{m}\coloneqq\{v\in\mathbb{R}^{m}:v_{m}\geq 0\}$. Let $L(\tilde{\gamma})$ be the Euclidean length of a curve $\tilde{\gamma}:[0,1]\to\mathbb{R}^{m}$.
@@
to see $\Psi$ as a function of the two last variables only, the two first being frozen. Observe that for any $(\xi,t,\zeta)$ in $[-1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)$,
@@
By compactness of $[-2,2]^{m-1}\times[0,\epsilon/2]$ and continuity of the maps $\Psi_{(\xi,t)}^{-1}$, we get that there exists a common $r_{0}>0$ such that the previous holds for any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$. Consider $r<r_{0}$ and $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, then
@@
×$\times\left(f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))\right)\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)$ (209)
@@
which shows that $G_{r}(x)=O(1)$ for $x$ such that $d(x,\partial M)<2r$.
@@ 6.2. Energy comparison
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\sqrt{\det(g_{x})\det(g_{y})}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
@@ 6.3. Proof of Theorem 3
We are now in a position to prove Theorem 3. Recall the context of this result : $(r_{n})\subset(0,+\infty)$ is a sequence such that $r_{n}\to 0$, $(M^{m},g)$ is a compact, connected, smooth Riemannian manifold with $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$ ), $k$ is a positive integer, $\mu_{k}$ is the $k$-th lowest Laplace (resp. Neumann) eigenvalue of $\Delta_{g}$, and $f_{k,r_{n}}$ is an eigenfunction of $-\tilde{\Delta}_{r_{n}}$ associated with the $k$-th eigenvalue $\lambda_{k}(-\tilde{\Delta}_{r_{n}})$ of this operator.

### v3

@@ 5. $L^{2}$ convergence
In this smooth context, the function $V(\cdot,r)$ is obviously continuous for any $r>0$. Since $M$ is compact, this implies that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies ($\mathrm{II}_{r}$) and ($\mathrm{I}_{r}$). Then Lemma 3.3 applies and ensures that $\tilde{\Delta}_{r}$ is a bounded self-adjoint operator acting on $L^{2}(M,\mathrm{vol}_{g})$. The compactness of $M$ also ensures that $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is locally Ahlfors regular: there exists a constant $C>1$ such that for any $x\in M$ and $r\in(0,\mathrm{diam}(M)]$,
@@ 5.1. Convergence in the sense of distributions
[…]
@@ 5.3. Uniform bound
We wish now to provide a uniform $L^{\infty}$ bound for the functions $\tilde{\Delta}_{r}\psi$ where $r$ is in a neighborhood of zero.
@@
for some positive definite, symmetric (m1)(m-1)-square matrix $A_{(\xi,t)}$. From the non-degeneracy of the metric and a Lipschitz bound, we have that there exist $C,\tilde{c}>0$ such that for all $t,s\in[0,\epsilon],\xi\in[-3,3]^{m-1},v\in\mathbb{R}^{m}$ we have
@@
[…]
@@
[…]
@@
to see $\Psi$ as a function of the two last variables only, the two first being frozen. Observe that for any $(\xi,t,\zeta)$ in $1,1]^{m-1}\times[0,\epsilon]\times\mathbb{B}^{m-1}_{\delta}(0)$,
@@
By compactness of $2,2]^{m-1}\times[0,\epsilon/2]$ and continuity of the maps $\Psi_{(\xi,t)}^{-1}$, we get that there exists a common $r_{0}>0$ such that the previous holds for any $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$. Consider $r<r_{0}$ and $(\xi,t)\in[-2,2]^{m-1}\times[0,\epsilon/2]$, then
@@
$\times\left(f(\Psi_{(\xi,t)}(\zeta,s))-f(\Psi_{(\xi,t)}(0,t))\right)\det\left(\left[\Psi_{(\xi,t)}^{*}g)\right]_{\zeta,s}\right)^{\frac{1}{2}}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(\zeta,s)$ (209)
@@
which shows that $G_{r}(x)=O(1)$ for $x$ such that $d(x,\partial M)<2r$. ∎
@@ 6.2. Energy comparison
$\qquad\qquad\qquad\qquad\qquad\qquad\times\sqrt{\det(g_{x})\det(g_{y})}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)\mathop{}\!\mathrm{d}\mathcal{L}^{m}(x)$
@@ 6.3. Proof of Theorem 3
We are now in a position to prove Theorem 3. Recall the context of this result : $(r_{n})\subset(0,+\infty)$ is a sequence such that $r_{n}\to 0$, $(M^{m},g)$ is a compact, connected, smooth Riemannian manifold with $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$ ), $k$ is a positive integer, $\mu_{k}$ is the $k$-th lowest Laplace (resp. Neumann) eigenvalue of $\Delta_{g}$, and $f_{k,r_{n}}$ is an eigenfunction of $\tilde{\Delta}_{r_{n}}$ associated with the $k$-th eigenvalue $\lambda_{k}(-\tilde{\Delta}_{r_{n}})$ of this operator.